cam0621:

The yearly cost in dollars, y, at a video game arcade based on total game tokens purchased, x, is y = y equals StartFraction 1 Over 10 EndFraction x plus 60.x + 60 for a member and y = y equals StartFraction 1 Over 5 EndFraction x. x for a nonmember. Explain how the graph of a nonmember’s yearly cost will differ from the graph of a member’s yearly cost.

Extrinix:

@cam0621 , can you send a picture of the question please?

AZ:

\( y = \dfrac{1}{10}x + 60\) for a member and \(y = \dfrac{1}{5}x\) for a nonmember

AZ:

Both equations are in the form of y = mx + b where m is the slope and b is the y-intercept

AZ:

@cam0621 Can you graph the two equations? You can use something like https://www.desmos.com/calculator to help you graph it so you can see how they look like

AZ:

So how will the graph for a nonmember look different from the graph of a member? Remember that y is the yearly cost which depends on how many game tokens or x that you buy.